# Understanding Transformers via $N$-gram Statistics

### v1

@@ Abstract
Transformer based large-language models (LLMs) display extreme proficiency with language yet a precise understanding of how they work remains elusive. One way of demystifying transformer predictions would be to describe how they depend on their context in terms of simple template functions. This paper takes a first step in this direction by considering families of functions (i.e. rules) formed out of simple $N$-gram based statistics of the training data. By studying how well these rulesets approximate transformer predictions, we obtain a variety of novel discoveries: a simple method to detect overfitting during training without using a holdout set, a quantitative measure of how transformers progress from learning simple to more complex statistical rules over the course of training, a model-variance criterion governing when transformer predictions tend to be described by $N$-gram rules, and insights into how well transformers can be approximated by $N$-gram rulesets in the limit where these rulesets become increasingly complex. In this latter direction, we find that for 78% of LLM next-token distributions on TinyStories, their top-1 predictions agree with those provided by our $N$-gram rulesets.
@@ 1 Introduction
    (Approximation-Variance Association) We observe an approximation-variance association indicating that next token LLM predictions that have low variance (across different training runs [^2] Different runs have different dataset shuffles.) tend to be well-approximated by $N$-gram rules. (Section 5)
@@
    (Approximation Strength) We study how well LLM predictions can be approximated by our $N$-gram rulesets, noting that significant gains in top1-accuracy occur as we increase ruleset complexity and diversity, whereby we achieve up to 78% top-1 accuracy on TinyStories (Table 2). We also visually ground these approximations with concrete examples (Figure 5), which may form the basis for dataset attribution methods in future work. (Section 7)
@@ 2 Related Work
Rule extraction methods for neural networks have been studied in quite different settings, e.g. (Jacobsson,, 2005; Mcmillan et al.,, 1991). Some recent works have performed $N$-gram analyses for large-language models in the setting of in-context learning (Akyürek et al.,, 2024) and associative recall (Arora et al.,, 2023). The “infini-gram" model (Liu et al.,, 2024) compares LLM predictions with the single $N$-gram rule given by retrieving the largest possible matching context from the training data. Our work uses shorter but more sophisticated $N$-gram rules. In (Voita et al.,, 2023), an approach to understanding how LLMs process $N$-grams is carried out at the level of individual neurons. This complements our dataset-based work, which treat models as a black box. In (Edelman et al.,, 2024), the evolution of the type of $N$-gram statistics that transformers learn during training is analyzed in the setting of synthetic Markov chain data, in contrast to our natural language setting. Other works studying the learning trajectory of language models include (Chen et al.,, 2024; Choshen et al.,, 2022). There is a large literature on building more sophisticated $N$-gram models, e.g. (Kneser and Ney,, 1995; Goodman,, 2001). Such models could have been incorporated into our set of rules, but for simplicity we choose not to include them.
@@ 3 Experimental Setup
We train standard decoder-only transformer models on the TinyStories (Eldan and Li,, 2023) dataset (480M tokens) consisting of children’s stories synthetically generated from GPT-4 using templated prompts. The value of this dataset lies in its linguistic simplicity, whereby it is possible to model language well on the dataset using very small models. Unless stated otherwise, our experiments use a 160M parameter model trained for 4 epochs, which achieves a loss of around 1.11 nats on the validation set. We train for 4 epochs since we use learning rate warmup and cosine learning rate decay and we want to ensure all datapoints receive updates with a high learning rate (this way all $N$-gram statistics have a fair chance of being learned during training). For overfitting experiments in Section 6.2, we train a 1.4B model for 10 epochs. In the Appendix, we include some additional corresponding experiments on Wikipedia (from MassiveText (Rae et al.,, 2022)) with a single epoch of training in order to validate that our results are of a general nature and extend to more complex datasets. For a fixed dataset, the only source of randomness among different runs are different dataset shuffles. Full experimental details are described in the Appendix.
@@ 4 $N$-Gram Rules
When $\sigma$ consists of all $+$ symbols, we get vanilla $N$-gram rules derived from the suffix of $C$. When $\sigma$ consists of $\pm$ symbols, we get vanilla $N$-gram rules derived from subsets of $C$. Varying the length and the entries of $\sigma$ yields the following rulesets888 There is some redundancy among the $\sigma$’s in terms of the rules they determine: for instance, in between any two $+$ consecutive symbols, permuting the order of -- and * will determine the same rule. Also in practice, we can assume the first entry of $\sigma$ is a $+$ since marginalizing the first token is equivalent to reducing the context length. From this, it follows that the number of distinct rules in $\mathcal{R}^{all}_{M}$ is $2,5,13,34,89,233,378$, for $M=1,\ldots,7$, respectively. :
@@ 5 Approximating Transformer Predictions with Rules
Let $p(t|C)$ denote the next-token distribution of an LLM conditioned on the context $C$ and for notational similarly, write $p_{r}(t|C)$ for $r(t|C)$, where $r$ is one of the rules defined in Section 4. We wish to measure how similar these distributions are (higher similarity corresponds to a better rule description). To that end, we use the variational distance to measure the difference of two distributions (we discuss our choice in the Appendix):
$d(p,q)=\frac{1}{2}\sum_{i}|p_{i}-q_{i}|.$ (10)
Since variational distance may be lacking in concrete interpretability, we will sometimes use top1-accuracy to measure similarity, defined by
@@
optimal rule distance: the minimum (possibly averaged over runs) distance between LLM predictions and rule predictions $\underset{r\in\mathcal{R}}{\min}\;\mathrm{avg}_{i}d(p_{i}(t|C),p_{r}(t|C))$
optimal rule: a rule achieving the optimal distance $\underset{r\in\mathcal{R}}{\mathrm{argmin}}\;\mathrm{avg}_{i}d(p_{i}(t|C),p_{r}(t|C))$
model variance: the average of the pairwise distance between LLM predictive distributions over different runs $\underset{i,j\atop\textrm{distinct runs}}{\mathrm{avg}}d(p_{i}(t|C),p_{j}(t|C))$
Table 1: Terminology associated to a context $C$. Here, $\mathcal{R}$ is some reference ruleset under consideration.
We are interested in determining the optimal rule $p_{r}(t|C)$ (as defined in Table 1) and if it has small optimal rule distance then we regard the rule as being a good description of the corresponding transformer prediction(s) $p_{i}(t|C)$. [^9] In practice, we will only have one model available and our optimal rules are computed per-context and per-model. In this section, we have available five models from five runs for use in computing optimal quantities. As a first step, note there is a distinguished rule
@@
Our main finding is an approximation-variance association: contexts with low model-variance tend to have low optimal rule distance. The surprising aspect of this association is the sufficiency of low model-variance (necessity is a given). [^12] Predictions which have high variance cannot be well approximated by a single model-independent rule. We use five runs in our analysis here since approximation by a rule that remains fixed across models yields a fortiori approximation by a per-model rule. We present the case of $7$-gram contexts in Figure 2 to corroborate this association, with additional examples relegated to the Appendix. We sample around six-thousand $7$-grams from the training data, sampling from logarithmically spaced buckets based on counts, and plot various relations between counts, model variances, and rule distances. For simplicitly, we consider the ruleset $\mathcal{R}=\mathcal{R}^{\textrm{suffix}}_{7}$ to limit the number of rules under consideration. Our analysis of Figure 2 can be summarized as follows:
Plot (d) summarizes our approximation-variance association. There is a clear positive correlation and reasonable fit between optimal rule distance vs model variance when using the ruleset $\mathcal{R}^{\textrm{suffix}}_{7}$. This fit weakens due to many outliers if we only consider the single $p_{\textrm{full}}$ rule (a vanilla $8$-gram model) as shown in (b). These outliers correspond to many LLM predictions being poorly approximated by $p_{\textrm{full}}$. The transition from (b) to (d) is a way of visualizing LLMs performing back-off, whereby LLMs rely on statistics from subsets of the context. We include plots (a) and (c) to highlight how replacing model-variance with the count of a context would lead to a much worse fit. We highlight count of the context $C$ because it is the most obvious and naive measure of how well one should expect $p(t|C)$ to match the rule $p_{\textrm{full}}(t|C)$. Nevertheless, the weak correlation between count and distance measures makes sense: an LLM can make predictions based on subcontexts of $C$ and those subcontexts can have very different count statistics than those of $C$.
@@
We believe our approximation-variance association and its corresponding analyses have significance beyond the experiments carried out here since they (i) highlight that naive count-based statistics do not provide the strongest signal in terms of how LLMs leverage dataset statistics (ii) suggest that LLM predictions that have low-variance are likely the ones that are amenable to description (or even explanation) by some underlying dataset statistic (with high-variance predictions being regarded as noise). We leave a more systematic exploration of (ii) to future work.
slope $=-0.20$ ; $R^{2}=0.11$
@@
Figure 2: TinyStories $7$-grams. Every point in the above plots represents a $7$-gram context. Shaded regions are plots obtained by bucketing along the x-axis and computing one standard deviation within the mean along the y-axis. Slope and $R^{2}$ values of plots are with respect to the linear fit of the data given by their axes. Optimal rule distances and model variances are computed with respect to five model runs. (a): $d(p(t|C),p_{\textrm{full}}(t|C))$ vs count of $C$. (b): $d(p(t|C),p_{\textrm{full}}(t|C))$ vs model variance. (c): model variance vs count of $C$. (d): similar to (b) but now the y-axis is optimal rule distance of the optimal rule from $\mathcal{R}^{\textrm{suffix}}_{7}$.
@@ 6.1 Curriculum Learning
Figure 3: Training Dynamics. Left: Models reach their lowest distance to more complex rules later in training. For rules with four tokens of context or less, the variational distance eventually starts increasing later in training. For six and seven tokens of context, the variational distance continues to decrease. Center & Right: The optimal rule selected always has nonincreasing distance and nondecreasing top1-accuracy relative to the ground truth (interpreted as a one-hot distribution), despite distances eventually increasing or plateauing for rules with less than six tokens of context. This shows that the optimal rule selection is improving with additional training even if the optimal rule distance with respect to model predictions is not improving. (One can imagine the rule predictions as a mesh in probability space, with LLM predictions navigating this space through training. The distance to the mesh may plateau but which rule is closest can continue to change.)
@@ 6.2 Early Stopping Criterion
Figure 4: Overfitting Detection. We plot both train loss (solid lines) and validation loss (dashed lines) for the full transformer and limited context length transformers (the latter are marked with an “x" for emphasis) on TinyStories. Unlike the full transformer which overfits, those with limited context length have train and validation loss curves closely following each other.
Figure 5: Rule selection for a TinyStories validation sequence. The above is a sequence from a heldout story. In the second and third columns are the ground truth, token by token, along with the rule context (as defined in Section 4) associated to the optimal rule from $\mathcal{R}_{7}^{all}$. The heatmap indicates the variational distance between optimal rule and LLM next token distributions at the given token. The first column shows at most two tokens, which are chosen as follows: If the LLM top-1 prediction disagrees with the ground truth, the LLM prediction is shown. If in addition, the rule selected makes a different top-1 prediction from the transformer, that token is also shown and the corresponding ground truth token is colored red. Thus red tokens are precisely the locations of disagreement between LLM and optimal rule greedy predictions. The last column shows the number of contexts supporting the optimal rule.
@@ 7 Rule Peformance
We see significant gains in accuracy at large $M$ when adding additional types of rules (for $M=7$ we gain $6\%$ each time in going from “suffix" to “subgram" to “all"). In the end, we are able to obtain 78% top1-accuracy between the per-prediction optimal rule and the LLM predictions, averaged over all tokens. This is perhaps a remarkably high figure, considering that the top1 accuracy of the model with respect to the ground truth on the validation set is 69.6%. At minimum, we have provided a precise quantification of structure in LLM next-token predictions: they are often matched (as measured by top token prediction) by some simple $N$-gram rule derived from the training data. See Section D.1 for some supplementary analysis.
@@
ruleset / $M$ 1 2 3 4 5 6 7
$\mathcal{R}^{\textrm{all}}_{M}$ 30.1 44.9 54.3 62.4 68.8 74.0 77.9
$\mathcal{R}^{\textrm{subgram}}_{M}$ 30.1 44.6 53.1 60.0 64.8 68.4 71.0
$\mathcal{R}^{\textrm{suffix}}_{M}$ 30.1 44.4 52.2 57.8 61.5 63.8 65.5
$\textrm{backoff}_{M}$ 30.1 42.5 48.7 52.6 54.6 55.8 56.6
Table 2: Approximation Strength. We look at the average top1-accuracy of the optimal rule versus LLM predictions for rules of varying strength and maximum context length. We compute this average over each token prediction from 100 random validation stories (around 22K tokens total).

### v2

@@ Abstract
Transformer based large-language models (LLMs) display extreme proficiency with language yet a precise understanding of how they work remains elusive. One way of demystifying transformer predictions would be to describe how they depend on their context in terms of simple template functions. This paper takes a first step in this direction by considering families of functions (i.e. rules) formed out of simple $N$-gram based statistics of the training data. By studying how well these rulesets approximate transformer predictions, we obtain a variety of novel discoveries: a simple method to detect overfitting during training without using a holdout set, a quantitative measure of how transformers progress from learning simple to more complex statistical rules over the course of training, a model-variance criterion governing when transformer predictions tend to be described by $N$-gram rules, and insights into how well transformers can be approximated by $N$-gram rulesets in the limit where these rulesets become increasingly complex. In this latter direction, we find that for 79% and 68% of LLM next-token distributions on TinyStories and Wikipedia, respectively, their top-1 predictions agree with those provided by our $N$-gram rulesets.
@@ 1 Introduction
    (Approximation Criterion) We observe that next token LLM predictions tend to be well-approximated by $N$-gram rules when the predictions have low variance across different training runs [^2] Different runs have different dataset shuffles.. In particular, this includes predictions conditioned on contexts with sufficiently high count in the training data. (Section 5)
@@
    (Approximation Strength) We study how well LLM predictions can be approximated by our $N$-gram rulesets, noting that significant gains in top1-accuracy occur as we increase ruleset complexity and diversity, whereby we achieve up to 79% top1-accuracy on TinyStories (Tables 2 and 5). We also visually ground these approximations with concrete examples (Figure 5), which may form the basis for dataset attribution methods in future work. Corresponding experiments on Wikipedia are shown in the Appendix. (Section 7)
We also open source our training datasets and related $N$-gram statistics so that others can verify and build upon our work. [^3] https://github.com/google-deepmind/transformer_ngrams
@@ 2 Related Work
Rule extraction methods for neural networks have been studied in quite different settings, e.g. (Jacobsson,, 2005; Mcmillan et al.,, 1991). Some recent works have performed $N$-gram analyses for large-language models in the setting of in-context learning (Akyürek et al.,, 2024; Olsson et al.,, 2022) and associative recall (Arora et al.,, 2023). The “infini-gram" model (Liu et al.,, 2024) compares LLM predictions with the single $N$-gram rule given by retrieving the largest possible matching context from the training data. Our work uses shorter but more sophisticated $N$-gram rules. In (Voita et al.,, 2023), an approach to understanding how LLMs process $N$-grams is carried out at the level of individual neurons. This complements our dataset-based work, which treat models as a black box. See also Svete and Cotterell, (2024) which studies how transformers can represent $N$-gram models. In (Edelman et al.,, 2024), the evolution of the type of $N$-gram statistics that transformers learn during training is analyzed in the setting of synthetic Markov chain data, in contrast to our natural language setting. Other works studying the learning trajectory of language models include (Chen et al.,, 2024; Choshen et al.,, 2022). There is a large literature on building more sophisticated $N$-gram models, e.g. (Kneser and Ney,, 1995; Goodman,, 2001). Such models could have been incorporated into our set of rules, but for simplicity we choose not to include them.
@@ 3 Experimental Setup
We train standard decoder-only transformer models on the TinyStories (Eldan and Li,, 2023) dataset (480M tokens) consisting of children’s stories synthetically generated from GPT-4 using templated prompts. The value of this dataset lies in its linguistic simplicity, whereby it is possible to model language well on the dataset using very small models. Unless stated otherwise, our experiments use a 160M parameter model trained for 4 epochs, which achieves a loss of around 1.11 nats on the validation set. We train for 4 epochs since we use learning rate warmup and cosine learning rate decay and we want to ensure all datapoints receive updates with a high learning rate (this way all $N$-gram statistics have a fair chance of being learned during training). For overfitting experiments in Section 6.2, we train a 1.4B model for 10 epochs. In the Appendix, we include additional corresponding experiments on Wikipedia (from MassiveText (Rae et al.,, 2022)) with a single epoch of training in order to validate that our results are of a general nature and extend to more complex datasets. For a fixed dataset, the only source of randomness among different runs are different dataset shuffles. Full experimental details are described in the Appendix.
@@ 4 $N$-Gram Rules
When $\sigma$ consists of all $+$ symbols, we get vanilla $N$-gram rules derived from the suffix of $C$. When $\sigma$ consists of $\pm$ symbols, we get vanilla $N$-gram rules derived from subsets of $C$. Varying the length and the entries of $\sigma$ yields the following rulesets999 There is some redundancy among the $\sigma$’s in terms of the rules they determine: for instance, in between any two $+$ consecutive symbols, permuting the order of -- and * will determine the same rule. Also in practice, we can assume the first entry of $\sigma$ is a $+$ since marginalizing the first token is equivalent to reducing the context length. From this, it follows that the number of distinct rules in $\mathcal{R}^{all}_{M}$ is $2,5,13,34,89,233,378$, for $M=1,\ldots,7$, respectively. :
@@ 5 Approximating Transformer Predictions with Rules
Let $p(t|C)$ denote the next-token distribution of an LLM conditioned on the context $C$ and for notational similarly, write $p_{r}(t|C)$ for $r(t|C)$, where $r$ is one of the rules defined in Section 4. We wish to measure how similar these distributions are (higher similarity corresponds to a better rule description). To that end, we use the variational distance to measure the difference of two distributions (we discuss our choice and others in the Appendix):
$d(p,q)=\frac{1}{2}\sum_{\alpha}|p_{\alpha}-q_{\alpha}|,$ (10)
where the summation is over the vocabulary index (i.e. the components of the probability vectors). Since variational distance may be lacking in concrete interpretability, we will sometimes use top1-accuracy to measure similarity, defined by
@@
Table 1: Terminology associated to a context $C$. Here $\mathcal{R}$ is some reference ruleset under consideration. The superscript on $p^{(i)}(t|C)$ is meant to denote the predictions of the $i$th model. In Section 5, we consider rules that are fixed across model runs (where we have five models) whereas elsewhere we will only have a single model (and thus optimal rules will be model specific).
optimal rule distance: the minimum (possibly averaged over runs) distance between LLM predictions and rule predictions $\underset{r\in\mathcal{R}}{\min}\;\mathrm{avg}_{i}d(p^{(i)}(t|C),p_{r}(t|C))$
optimal rule: a rule achieving the optimal distance $\underset{r\in\mathcal{R}}{\mathrm{argmin}}\;\mathrm{avg}_{i}d(p^{(i)}(t|C),p_{r}(t|C))$
model variance: the average of the pairwise distance between LLM predictive distributions over different runs $\underset{i,j\atop\textrm{distinct runs}}{\mathrm{avg}}d(p^{(i)}(t|C),p^{(j)}(t|C))$
We are interested in determining the optimal rule $p_{r}(t|C)$ (as defined in Table 1) and if it has small optimal rule distance then we regard the rule as being a good description of the corresponding transformer predictions. [^10] In practice, we will only have one model available and our optimal rules are computed per-context and per-model. In this section, we have available five models from five runs for use in computing optimal quantities. As a first step, note there is a distinguished rule
@@
One of our main findings is an approximation criterion: contexts that have low model-variance tend to have low optimal rule distance. In particular, this includes contexts with sufficiently high frequency in the training data. The latter situation is to be expected: the more often a context $C$ occurs in the training data, the more the minimization of the cross entropy loss objective encourages the network to make predictions close to $p_{\textrm{full}}(t|C)$.
@@
The novel aspect of our approximation criterion is the sufficiency of low model-variance situation even in cases when the context is rare. [^13] Necessity is a given. Predictions which have high variance cannot be well approximated by a single model-independent rule. We use five runs in our analysis here since approximation by a rule that remains fixed across models yields a fortiori approximation by a per-model rule. We present the case of $7$-gram contexts in Figure 2 to corroborate the approximation criterion, with additional examples relegated to the Appendix. We sample around six-thousand $7$-grams from the training data, sampling from logarithmically spaced buckets based on counts, and plot various relations between counts, model variances, and rule distances. For simplicitly, we consider the ruleset $\mathcal{R}=\mathcal{R}^{\textrm{suffix}}_{7}$ to limit the number of rules under consideration. Our analysis of Figure 2 can be summarized as follows:
Plot (a) shows how with increasing count of the number occurrences of the context $C$ in the training data, LLM predictions become nearer to $p_{\textrm{full}}(t|C)$, which in this case, is the vanilla $8$-gram rule. Nevertheless, for all but the highest of counts, we have a large spread of distances: even for unique $7$-gram contexts, some predictions are well-approximated by $p_{\textrm{full}}(t|C)$ while others are close to having disjoint-support (distance equal to $1$). Plot (c) also shows that while model variance decreases with count of the context (as expected) we have a large spread of model variances for contexts with intermediate or low counts. Since the contexts whose predictions have high model variance can be regarded as “noise", one can ask whether those contexts with low model variance have some structure. Plots (b) and (c) address this question. While for (b), we see that the 8gram-rule has a large spread when plotted against model variance, there is a significant reduction in outliers in (d) when the y-axis is the optimal distance to rules in $\mathcal{R}^{\textrm{suffix}}_{7}$. Concretely, the transition from (b) to (d) is a way of visualizing LLMs performing back-off, whereby LLMs rely on statistics from subsets of the context. Moreover, the lower left portion of (d) is a manifestation of our approximation criterion: contexts that yield consistent predictions across model runs (i.e. sufficiently low model variance) are indicative of rule-like behavior (in this case, good approximation with a suffix $N$-gram rule formed out of the context).
@@
We believe our approximation criterion and its corresponding analyses have significance beyond the experiments carried out here since they (i) highlight that naive count-based statistics do not provide the strongest signal in terms of how LLMs leverage dataset statistics (since as Figure 2(a) shows, high count can still have high model variance) (ii) suggest that LLM predictions that have low-variance are likely the ones that are amenable to description (or even explanation) by some underlying dataset statistic (with high-variance predictions being regarded as noise). We leave a more systematic exploration of (ii) to future work.
slope $=-0.05$ ; $R^{2}=0.35$
@@
Figure 2: TinyStories $7$-grams. Every point in the above plots represents a $7$-gram context. Shaded regions are obtained by bucketing along the x-axis and computing one standard deviation within the mean along the y-axis. Slope and $R^{2}$ values of plots are with respect to the linear fit of the data. Optimal rule distances and model variances are computed with respect to five model runs. (a): $d(p(t|C),p_{\textrm{full}}(t|C))$ vs count of $C$. (b): $d(p(t|C),p_{\textrm{full}}(t|C))$ vs model variance. (c): model variance vs count of $C$. (d): similar to (b) but now the y-axis is optimal rule distance of the optimal rule from $\mathcal{R}^{\textrm{suffix}}_{7}$. Model size: 160M.
@@ 6.1 Curriculum Learning
Figure 3: Training Dynamics. Left: Models reach their lowest distance to more complex rules later in training. For rules with four tokens of context or less, the variational distance eventually starts increasing later in training. For six and seven tokens of context, the variational distance continues to decrease. Center & Right: The optimal rule selected always has nonincreasing distance and nondecreasing top1-accuracy relative to the ground truth (interpreted as a one-hot distribution $p_{\mathrm{gt}}$), despite distances to model predictions eventually increasing or plateauing for rules with less than six tokens of context. This shows that the optimal rule selection is improving with additional training even if the optimal rule distance with respect to model predictions is not improving. (One can imagine the rule predictions as a mesh in probability space, with LLM predictions navigating this space through training. The distance to the mesh may plateau but which rule is closest can continue to change.) Model size: 160M.
@@ 6.2 Early Stopping Criterion
Figure 4: Overfitting Detection. We plot both train loss (solid lines) and validation loss (dashed lines) for the full transformer and limited context length transformers (the latter are marked with an “x" for emphasis) on TinyStories. Unlike the full transformer which overfits, those with limited context length have train and validation loss curves closely following each other. Model size: 1.4B.
Figure 5: Rule selection for a TinyStories validation sequence. The above is a sequence from a heldout story. In the second and third columns are the ground truth, token by token, along with the rule context (as defined in Section 4) associated to the optimal rule from $\mathcal{R}_{7}^{\textrm{all}}$. The heatmap indicates the variational distance between optimal rule and LLM next token distributions at the given token. The first column shows at most two tokens, which are chosen as follows: If the LLM top-1 prediction disagrees with the ground truth, the LLM prediction is shown. If in addition, the rule selected makes a different top-1 prediction from the transformer, that token is shown as the second token and the corresponding ground truth token is colored red. Thus red tokens are precisely the locations of disagreement between LLM and optimal rule greedy predictions. The last column shows the number of contexts supporting the optimal rule. Model size: 160M.
@@ 7 Rule Peformance
We see significant gains in accuracy at large $M$ when adding additional types of rules. In the end, we are able to obtain 78% top1-accuracy between the per-prediction optimal rule and the LLM predictions, averaged over all tokens. This is perhaps a remarkably high figure, considering that the top1 accuracy of the model with respect to the ground truth on the validation set is 69.6%. At minimum, we have provided a precise quantification of structure in LLM next-token predictions: they are often matched (as measured by top token prediction) by some simple $N$-gram rule derived from the training data. See Section D.1 for some supplementary analysis. Using the $L^{\infty}$ distance instead of the variational distance gives us a slightly higher result of 79%, see Table 5.
@@
Table 2: Top-1 accuracy of optimal rule. We look at the average top1-accuracy of the optimal rule versus LLM predictions for rules of varying strength and maximum rule context length $M$. We compute this average over each token prediction from 100 random validation stories (around 22K tokens total). Model size: 160M.
ruleset / context length 1 2 3 4 5 6 7
$\mathcal{R}^{\textrm{all}}_{M}$ 30.0 44.8 54.3 62.4 68.8 74.0 78.0
$\mathcal{R}^{\textrm{subgram}}_{M}$ 30.0 44.5 53.1 60.0 64.8 68.5 71.1
$\mathcal{R}^{\textrm{suffix}}_{M}$ 30.0 44.4 52.2 57.9 61.5 63.8 65.6
$\textrm{backoff}_{M}$ 30.0 42.5 48.7 52.7 54.6 55.9 56.7